OpenStudy (anonymous):

(a + b - c)(a + b + c) I'm getting a^2+b^2-c^2+2bc but this isn't right. Little help?

OpenStudy (skullpatrol):

Use the distributive property.

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (kc_kennylau):

I would group \(a+b\) together and use \((\alpha+\beta)(\alpha-\beta)=\alpha^2-\beta^2\) :)

OpenStudy (skullpatrol):

$$(a + b - c)(a + b + c) $$ $$ =(a + b - c)a + (a + b - c)b + (a + b - c)c$$

OpenStudy (skullpatrol):

Now use this form of the distributive property:$$\Huge (b+c)a=ba+ca$$

OpenStudy (anonymous):

@skullpatrol I think kc-kennylau gave us the easiest way.

OpenStudy (skullpatrol):

Let the student decide please :-)

OpenStudy (anonymous):

I am sorry.

OpenStudy (skullpatrol):

np

OpenStudy (anonymous):

Figured it out! thanks a bunch for your help (:

OpenStudy (anonymous):

Took me a few minutes but i got it! :)

OpenStudy (skullpatrol):

Thanks for trying to understand :)

OpenStudy (anonymous):

Thank-you for your time!

OpenStudy (skullpatrol):

:-)